OpenStudy (anonymous):

What is the value of y in the equation 2(3y - 4) = 10 ? Please explain how to get answer thanks!!

OpenStudy (anonymous):

2(3y - 4) = 10 6y - 8 = 10 6y = 18 y = 3

OpenStudy (anonymous):

3y-4=5 3y=9 y=3

OpenStudy (anonymous):

pk, you are too quick on that keyboard

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (anonymous):

Thanks

OpenStudy (curry):

2(3y - 4) = 10 6y - 8 = 10 6y = 18 y = 3